# Atomic Fact Decomposition Helps Attributed Question Answering

### v1

@@ Abstract.
Attributed Question Answer, Information Retrieval, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; April 28 - May 2; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06submissionid: 184ccs: Information systems Information retrievalccs: Computing methodologies Natural language generation
1. Introduction
Large Language Models (LLMs), pre-trained on large-scale text corpora (10387715, ), have demonstrated remarkable capabilities in natural language understanding and generation tasks (zhuang2023toolqa, ; wang2024large, ). However, despite their impressive performance, they often face significant challenges in real-world applications due to issues like lack of interpretability and indecisiveness (huang2023survey, ; peskoff-stewart-2023-credible, ). These limitations undermine the reliability and trustworthiness of LLMs, particularly in contexts where transparency and accuracy are essential. To address these issues, Attributed Question Answering (AQA) (li2023survey, ) has emerged as a solution, focusing on linking their generated answers to specific sources of evidence.
Figure 1. The motivation of previous methods and our proposed ARE. Previous methods directly use LLMs to generate sub-questions for the long-form answer and edit the whole answer. Our proposed ARE leverages instruction-tuned LLMs to achieve molecular-to-atomic stage, and backtrack the edited atomic facts to achieve atomic-to-molecular stage.
@@ 1. Introduction
Depending on the timing of retrieval, previous research can be divided into two categories: Retrieval-then-Read (RTR)  (gao2023enabling, ), and Post-hoc Retrieval (bohnet2022attributed, ; gao2023rarr, ; chen2023purr, ). RTR delivers relevant answers by retrieving documents based on the query, providing detailed context that can enhance the richness of the response. However, LLMs often link content to irrelevant or incorrect sources and lack an effective mechanism for subsequent correction. In contrast, post-hoc retrieval typically retrieves specific external information for initially generated long-form answers, effectively reducing linking errors and enabling targeted revisions, which enhances flexibility and robustness. (dhuliawala2023chainofverification, ; kim2024re, ).
@@
Despite the flexibility post-hoc retrieval offers in verifying each fact and correcting it in long-form answers, it introduces several significant challenges. Firstly, existing methods (chen2023purr, ; kim2024re, ) prompt LLMs to directly generate a series of sub-questions for retrieval based on the long-form answer (As shown in Figure 1 (I)). However, the complex logical structures and vast amounts of information typically found in long-form answers make it difficult for these methods to comprehensively capture all necessary facts. As a result, the generated sub-questions often fail to align with the core content of the answer, leading to irrelevant questions and ultimately hindering the retrieval of accurate supporting evidence. tribute methods always overlooking the broader context, which can lead to inaccurate or incomplete attribution.
@@
Moreover, existing methods (gao2023rarr, ; wu2024updating, ) that directly edit long-form answers holistically often face difficulty in precisely locating specific content for revision, making it challenging to preserve the original intent while ensuring consistency between the answer and the attribution. This limitation arises from either insufficient editing, where critical details remain inadequately revised, or excessive editing, which distorts the original intent or introduces unnecessary changes, which disrupts the coherence of the answer. Thus, effectively balancing the challenges of insufficient and excessive editing has become a critical issue.
@@
To address the above limitations, we propose a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for AQA (As shown is Figure 1 (II)). Specifically, ARE first prompts LLMs to generate long-form answers for the given question. Subsequently, ARE utilizes an instruction-tuned LLM, trained on a well-constructed fact decomposition dataset, to decompose the long-form answers into molecular clauses and then into atomic facts. These atomic facts are then used to search the evidences from a search engine.
@@
Figure 2. An example of the ARE process. In Molecular-to-Atomic stage, an instruction-tuned fact decomposition LLM decomposes the long-form answer $X$ into molecular clauses and atomic facts. Atomic facts are then used to retrieve evidences. In Atomic-to-Molecular stage, a Verifier assesses the relationship between the evidence and facts, resulting in three states: 1. supportive, requiring no further action; 2. editing required, necessitating revision to the atomic facts; and 3. irrelevant, requiring a new evidence retrieval. Finally, the atomic facts are backtracked to the original position of $X$ to generate the revised answer $X^{{}^{\prime}}$. The Attribution Report consists of all evidences $\{E_{1},...,E_{m}\}$, where $MF_{i}^{{}^{\prime}}$ and $e_{ij}^{{}^{\prime}}$ are the edited contents.
2. Related Work
2.1. Attributed Question Answering
@@
From the perspective of retrieval timing, two notable trends have recently emerged: (1) Retrieval-then-Read (RTR) and (2) Post-hoc Retrieval. Trivedi et al. (trivedi2022interleaving, ) introduce IRCoT, a method that interleaves retrieval with steps in a Chain of Thought (CoT). This approach not only guides the retrieval process using CoT but also utilizes the retrieved results to enhance the CoT.
@@
Muller et al. (muller-etal-2023-evaluating, ) investigate attribution in cross-lingual question answering (QA). ALCE (gao2023enabling, ) employs various methods to integrate retrieved documents into large language models (LLMs) for answer generation. Li et al. (li2023llatrieval, ) introduce a progressive selection of evidence using LLMs with a classification-based prompting template. Bohnet et al. (bohnet2022attributed, ) conduct an extensive evaluation of LLM attributions, finding that while the Retrieval-then-Read (RTR) approach performs well, it necessitates the comprehensive use of a traditional training set, thereby highlighting the potential of post-hoc retrieval. RARR (gao2023rarr, ) is the first framework to implement question decomposition-based retrieval followed by revision. Building on RARR, PURR presents an end-to-end editor for text revision (chen2023purr, ). Kang et al. (kang2023mitigating, ) propose a combination of RTR and post-hoc retrieval strategies. The limitations of question decomposition were discussed in the Introduction. Our approach addresses these limitations through molecular-to-atomic fact decomposition and atomic-to-molecular editing, resulting in significant performance improvements.
2.2. Fact Decomposition
@@
The technology of decomposition has been shown to effectively address complex questions, particularly in various reasoning tasks (lin2023decomposing, ; Ye_2023, ; zhang2023reasoning, ) and claim verification tasks (lin-etal-2023-decomposing, ; chen2023complex, ; kim2023kggpt, ; wang2023hallucination, ; kamoi-etal-2023-wice, ). While fact decomposition can accurately generate sub-facts that represent the original answer, it struggles to aggregate evidence for sentences based on the decomposition results. This limitation becomes evident when retrieving evidence by fact to support each sentence that contains more than one fact.
2.3. Hallucination Detection and Revision
Hallucination detection (chen2023hallucination, ; wang2023hallucination, ; kadavath2022language, ; mndler2023selfcontradictory, ; mishra2024fine, ) is a challenging yet essential task for improving the reliability of large language models (LLMs) in real-world scenarios. Similarly, Zheng et al. (ZLLP2024, ) introduced TrustScore, the first effective evaluation metric designed to assess the trustworthiness of LLM responses in a reference-free context.
@@ 2.3. Hallucination Detection and Revision
To tackle hallucinations, numerous model editing methods have been developed that do not involve updating the parameters of large language models (LLMs) (zheng-etal-2023-edit, ; gao2023rarr, ; HLTW+2023, ; dhuliawala2023chainofverification, ; wu2024updating, ; song2024knowledge, ). In contrast to our approach, most of these methods focus exclusively on either hallucination detection or revision.
3. Preliminary
@@
where $X$ represents the long-form answer, $A$ is an attribution report consisting of of a collection of evidence, and $\mathcal{M}_{\text{AQA}}$ is a model used for the AQA task.
@@
Symbol definitions. The key symbols used in our framework are depicted as follows: $MF_{i}$ represents a molecular clause within the long-form answer $X$; $AF_{ij}$ denotes the atomic fact of $MF_{i}$; $e_{ij}$ corresponds to the evidence for each atomic fact $AF_{ij}$; $E_{i}$ represents the aggregated evidence for the molecular clause $MF_{i}$, compiled from multiple evidence $e_{ij}$.
Entropy measures the uncertainty or information content associated with random variables. The higher the entropy, the more uncertain or random the information is (ash2012information, ). It can be calculated as:
@@
Hypothesis. Atomic facts exhibit lower entropy than sub-questions, leading to improved evidence retrieval. As demonstrated by Passalis et al. (7439840, ) that minimizing entropy can enhance performance in retrieval tasks. Long-form answers, characterized by greater complexity, have higher entropy, which makes it more challenging to retrieve relevant evidences. Traditional post-hoc retrieval based AQA methods use LLMs to generate sub-questions directly, which may contain multiple meanings and uncertainties. Thus, these sub-questions often have high entropy, leading to retrieving ambiguous or irrelevant evidence (As shown in Figure 1 (I)). In contrast, fact decomposition preserves higher certainty and lowers entropy, thereby enhancing evidence retrieval (ash2012information, ). The experimental results in Section 5.5 verify our hypothesis by demonstrating that atomic-fact based retrieval outperforms sub-question based retrieval.
@@
Figure 3. The process of dataset construction and instruction tuning. We first extract the neighbors of the corresponding entity from Wikidata, and then utilize an LLM to convert each triple into a single atomic fact. Subsequently, all facts are transformed into a text comprising multiple molecular clauses by the LLM. The atomic facts are then used to construct each sentence. After two prompts, we obtain the dataset for instruction-tuning with LoRA.
4. Approach
@@
In this section, a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework is proposed for AQA task. Specifically, our ARE framework contains three stages: 1) Molecular-to-Atomic Fact Decomposition generally decomposes the generated long-form answer into molecular clauses and atomic facts by an instruction-tuned LLM; 2) Atomic Fact based Evidence Retrieval employs a search engine to retrieve evidence and selects the most similar evidence using a reranking model; 3) LLM-based Evidence Verifing and Editing guides LLMs through a verified result to either expand atomic facts for further re-retrieval or edit atomic facts with retrieved evidence.
4.1. Molecular-to-Atomic Fact Decomposition
@@
4.1.1. Long-form answer generation
LLMs are trained on the large-scale corpora which endow them with powerful generative capabilities. Long-form answers are generated by feeding few-shot prompting template and a specific question to LLMs. Formally, the whole process can be represented as:
@@ 4.1.1. Long-form answer generation
4.1.2. Fact decomposition
Fact decomposition aims to decompose long-form answers characterized by complex logical structures into molecular clauses and multiple atomic facts. Specifically, our ARE framework constructs a molecular-to-atomic fact decomposition dataset for instruction-tuning LLM, and this process enhances LLM’s ability to effectively manage the intricacies of long-form answers.
Dataset Construction. To fine-tune the molecule-to-atomic fact decomposition LLM, ARE constructs an instruct-tuning dataset that converts triples into texts containing molecule clauses and atomic facts derived from Knowledge Graphs (KGs) using LLMs. As depicted in Figure 3, the dataset construction process is divided into two steps: data collection and data processing with LLMs.
Step1: ARE first randomly selects the entities from the Knowledge Graph Question Answering (KGQA) dataset (gu2021beyond, ) and then extracts the triplets of given entities from Wikidata [^1] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=. As some properties or objects in the triplets may be useless (e.g., JPEG, code, and ISSN), they are thus further filtered by heuristic rules in the extracting process. Details of the rules can be found in Appendix C.4.
Step2: ARE utilizes LLMs to randomly transform several triplets into atomic facts $\{AF_{i1},...,AF_{ik}\}$. The atomic facts are fed into LLMs to generate a response in the JSON format: {“Generated content”: $S$, “Molecular clauses”: $MF_{i}$, “Atomic facts”: $AF_{ij}$}, where $S$ is text consisting of $n$ molecular clauses, $MF_{i}$ is $i$-th molecular clause, $i\in\{1,...,m\}$, and $AF_{ij}$ is the $j$-th atomic fact of the $MF_{i}$, $j\in\{1,...,k\}$.
Through the above process, an instruction-tuning dataset can be obtained, which contains 7138 samples for training and 1000 for evaluation. An example data can be found in Appendix A.2.
Instruction-tuning the Molecule-to-Atomic LLM. For the fine-tuning stage, ARE employs the Low-Rank Adaptation (LoRA) technique (hu2022lora, ) for instruction-tuning the Llama3-8B-Instruct LLM to reduce computational complexity and time consumption. Specifically, LoRA works by freezing parameters $\theta_{0}$ of the pre-trained model, while adding trainable parameters $\Delta\theta_{0}$ that can be expressed as the product of two low-rank matrices:
(1) $\Delta\theta_{0}=\mathbf{BA},$

### v2

@@ Abstract
Attributed Question Answer, Information Retrieval, Large Language Models.
I Introduction
Large Language Models (LLMs), pre-trained on large-scale text corpora [1], have demonstrated remarkable capabilities in natural language understanding and generation tasks [2, 3]. However, their application in real-world scenarios is hindered by a lack of transparency, limited interpretability, and occasional factual inaccuracies, [4, 5].These limitations compromise the reliability of LLM outputs, especially in settings where factual correctness and evidence traceability are critical.
To address these challenges, Attributed Question Answering (AQA) [6] has emerged as a promising paradigm. AQA aims to enhance answer trustworthiness by explicitly linking generated responses to verifiable sources of evidence. As illustrated in Figure 1, when asked “Who has the most Super Bowl rings?”, an LLM initially responds: “The player with the most Super Bowl rings is Tom Brady. Tom Brady is an American football quarterback who has won six Super Bowl championships.” Upon retrieving external evidence, however, the system identifies a factual error — Tom Brady has won seven, not six, championships. This triggers a post-hoc correction process, resulting in a revised and properly attributed answer, along with the supporting evidence.
@@ I Introduction
Depending on the timing of retrieval, previous research can be divided into two categories: Retrieval-then-Read (RTR)  [7], and Post-hoc Retrieval [8, 9, 10, 11]. RTR delivers relevant answers by retrieving documents based on the query, providing detailed context that can enhance the richness of the response. However, LLMs often link content to irrelevant or incorrect sources and lack an effective mechanism for subsequent correction. In contrast, post-hoc retrieval typically retrieves specific external information for initially generated long-form answers, effectively reducing linking errors and enabling targeted revisions, which enhances flexibility and robustness [12, 13].
@@
Despite the flexibility post-hoc retrieval offers in verifying each fact and correcting it in long-form answers, it introduces several significant challenges. Firstly, existing methods [10, 13] prompt LLMs to directly generate a series of sub-questions for retrieval based on the long-form answer (As shown in Figure 1 (I)). However, the complex logical structures and vast amounts of information typically found in long-form answers make it difficult for these methods to capture all necessary facts comprehensively. As a result, the generated sub-questions often fail to align with the core content of the answer, leading to irrelevant questions and ultimately hindering the retrieval of accurate supporting evidence.
@@
Moreover, existing methods [9, 14] that directly edit long-form answers holistically often face difficulty in precisely locating specific content for revision, making it challenging to preserve the original intent while ensuring consistency between the answer and the attribution. This limitation arises from either insufficient editing, where critical details remain inadequately revised, or excessive editing, which distorts the original intent or introduces unnecessary changes, which disrupts the coherence of the answer. Thus, effectively balancing the challenges of insufficient and excessive editing has become a critical issue.
Figure 1: The motivation of previous methods and our proposed ARE. Previous methods directly use LLMs to generate sub-questions for the long-form answer and edit the whole answer. Our proposed ARE leverages instruction-tuned LLMs to achieve molecular-to-atomic stage, and backtrack the edited atomic facts to achieve atomic-to-molecular stage.
@@
To address the above limitations, we propose a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework for AQA (As shown is Figure 1 (II)). Specifically, ARE first prompts LLMs to generate long-form answers for the given question. Subsequently, ARE utilizes an instruction-tuned LLM [^1] The fine-tuned atomic fact decomposition LLM can be found in https://github.com/zhichaoyan11/FIDES, trained on a well-constructed fact decomposition dataset, to decompose the long-form answers into molecular clauses and then into atomic facts. These atomic facts are then used to search the evidence from a search engine.
@@
Figure 2: An example of the ARE process. In Molecular-to-Atomic stage, an instruction-tuned fact decomposition LLM decomposes the long-form answer $X$ into molecular clauses and atomic facts. Atomic facts are then used to retrieve evidences. In Atomic-to-Molecular stage, a Verifier assesses the relationship between the evidence and facts, resulting in three states: 1. supportive, requiring no further action; 2. editing required, necessitating revision to the atomic facts; and 3. irrelevant, requiring a new evidence retrieval. Finally, the atomic facts are backtracked to the original position of $X$ to generate the revised answer $X^{{}^{\prime}}$. The Attribution Report consists of all evidences $\{E_{1},...,E_{m}\}$, where $MF_{i}^{{}^{\prime}}$ and $e_{ij}^{{}^{\prime}}$ are the edited contents.
II Related Work
II-A Attributed Question Answering
@@
From the perspective of retrieval timing, two notable trends have recently emerged: (1) Retrieval-then-Read (RTR) and (2) Post-hoc Retrieval. Trivedi et al. [15] introduces IRCoT, a method that interleaves retrieval with steps in a Chain of Thought (CoT). This approach not only guides the retrieval process using CoT but also utilizes the retrieved results to enhance the CoT.
@@
Muller et al. [16] investigates attribution in cross-lingual question answering. ALCE [7] employs various methods to integrate retrieved documents into LLMs for answer generation. Li et al. [17] introduces a progressive selection of evidence using LLMs with a classification-based prompting template. Bohnet et al. [8] conducts an extensive evaluation of LLM attributions, finding that while the Retrieval-then-Read (RTR) approach performs well, it necessitates the comprehensive use of a traditional training set, thereby highlighting the potential of post-hoc retrieval. RARR [9] is the first framework to implement question decomposition-based retrieval followed by revision. Building on RARR, PURR presents an end-to-end editor for text revision [10]. Kang et al. [18] proposes a combination of RTR and post-hoc retrieval strategies. The limitations of question decomposition were discussed in the Introduction. Our approach addresses these limitations through molecular-to-atomic fact decomposition and atomic-to-molecular editing, resulting in significant performance improvements.
II-B Question or Fact Decomposition
@@
The technology of decomposition has been shown to effectively address complex questions, particularly in various Question Answering (QA) tasks (e.g., Multi Hop Question Answering (MHQA)) [19, 20, 21, 22] and claim verification tasks [23, 24, 25, 26, 27]. Under the MHQA task, [28, 29, 30, 31, 32] typically decompose the original question or introduce retrieval augment generation methods to mitigate the ambiguity of questions and the model parameter knowledge limitations, respectively. While fact decomposition can generate sub-facts that represent the original answer, it struggles to decompose long-form answers that have a complex semantic structure and aggregate evidence for sentences based on the decomposition results. This limitation becomes evident when retrieving evidence by fact to support each sentence that contains more than one fact.
II-C Hallucination Detection and Revision
Hallucination detection [33, 26, 34, 35, 36] is a challenging yet essential task for improving the reliability of LLMs in real-world scenarios. Similarly, Zheng et al. [37] introduced TrustScore, the first effective evaluation metric designed to assess the trustworthiness of LLMs responses in a reference-free context. Yan et al. [38] proposed CRAG, which trains an evidence evaluator model to assess the quality of retrieved evidence and classify retrieval actions as “Correct,” “Incorrect,” or “Ambiguous”. Different from our framework, CRAG aims to improve retrieval robustness by evaluating and refining retrieved evidence before generation and focuses on improving evidence quality rather than ensuring attribution and alignment with source material.
@@ II-C Hallucination Detection and Revision
To tackle hallucinations, numerous model editing methods have been developed that do not involve updating the parameters of LLMs [39, 9, 40, 12, 14, 41]. In contrast to our approach, most of these methods focus exclusively on either hallucination detection or revision.
III Preliminary
@@
where $X$ represents the long-form answer, $A$ is an attribution report consisting of a collection of evidence, and $\mathcal{M}_{\text{AQA}}$ is a model used for the AQA task.
@@
Symbol definitions. The key symbols used in our framework are defined as follows: $MF_{i}$ denotes a molecular clause in the long-form answer $X$ that contains at least one fact subject to further decomposition; $AF_{ij}$ represents an atomic fact derived from $MF_{i}$ that cannot be further decomposed; $e_{ij}$ refers to the evidence retrieved for each atomic fact $AF_{ij}$; and $E_{i}$ denotes the aggregated evidence for the molecular clause $MF_{i}$, compiled from the corresponding evidence pieces $e_{ij}$.
Entropy measures the uncertainty or information content associated with random variables. The higher the entropy, the more uncertain or random the information is [42]. It can be calculated as:
@@
Hypothesis. Atomic facts exhibit lower entropy than sub-questions, leading to improved evidence retrieval. As demonstrated by Passalis et al. [43] that minimizing entropy can enhance performance in retrieval tasks. Long-form answers, characterized by greater complexity, have higher entropy, which makes it more challenging to retrieve relevant evidences. Traditional post-hoc retrieval based AQA methods use LLMs to generate sub-questions directly, which may contain multiple meanings, uncertainties and lose some key entities. Thus, these sub-questions often have high entropy, leading to retrieving ambiguous or irrelevant evidence (As shown in Figure 1 (I)). In contrast, fact decomposition preserves higher certainty and lowers entropy, thereby enhancing evidence retrieval [42].
To support this hypothesis, we begin with the following assumptions:
1) $X$ is the long-form answer, $\mathrm{h}$ is a deterministic extractor that maps each $X$ to a set of entities: $E=\mathrm{h}(X)$, hence by the definition of conditional entropy for a deterministic function, $H(E|X)=0$.
2) The support set for atomic fact $AF$ is $S_{AF}(X)=\{af:P(AF=af|X)>0\}$ and for sub-question $S_{Q}(X)={q:P(Q=q|X)>0}$. To fully cover the same triple fact $(e_{h},r,e_{t})$, the LLMs must generate at least two sub-questions (one for head entity and one for tail). Hence
$\mid S_{Q}(X)\mid\geq 2\mid S_{AF}(X)\mid$ (1)
3) From atomic fact set $AF$, $E$ can be recovered, the entropy $H(E|AF)=0$, similarly, from sub-question set $Q$, the entropy $H(E|Q)=0$.
The theory of maximum entropy: For a length of support set $\mid\mathrm{supp(P)\mid}=M$, the upper bound of entropy is $H(P)\leq\mathrm{log}M$. The generation process of the support set is considered without any bias, the maximum entropy is uniform distribution. Therefore, we adopt $H(AF|X)\approx\mathrm{log}\mid S_{AF}(X)\mid$ and $H(Q|X)\approx\mathrm{log}\mid S_{Q}(X)\mid$.
From eq 1,
$\mathrm{log}\mid S_{Q}\mid\geq\mathrm{log}(2\mid S_{AF}\mid)=\mathrm{log}2+\mathrm{log}\mid S_{AF}\mid,$ (2)
consider the upper bound of entropy,
$H(Q|X)\geq\mathrm{log}2+\mathrm{log}\mid S_{AF}(X)\mid\geq\mathrm{log}2+H(AF|X)$ (3)
$H(Q|X)-H(AF|X)\geq\mathrm{log}2>0$ (4)
$\Rightarrow H(AF|X)<H(Q|X).$
The theory of chain rule of entropy: Given the two random variables $(A|B)$, the chain is:
$H(A,B)=H(A)+H(B|A)=H(B)+H(A|B).$ (5)
$\mathrm{For}\ AF\mathrm{:}\ H(E,AF)$ $=H(AF)+H(E|AF)$
$=H(AF)+0$
$=H(E)+H(AF|E)$
$\Rightarrow H(AF)=H(E)+H(AF|E).$
For $Q$,
$H(E,Q)$ $=H(Q)+H(E|Q)$
$=H(Q)+0$
$=H(E)+H(Q|E)$
$\Rightarrow H(Q)=H(E)+H(Q|E).$
Due to $H(E|X)=0$, $H(AF)=H(E)+H(AF|X)$, $H(Q)=H(E)+H(Q|X)$. The relation of $H(Q)$ and $H(AF)$ is
$H(Q)-H(AF)$
$=[H(E)+H(Q|X)]-[H(E)+H(AF|X)]$
$=H(Q|X)-H(AF|X)\geq\mathrm{log}2\Rightarrow H(Q)>H(AF).$
The entropy of atomic fact is lower than sub-question. The experimental results in Section V-H also verify our hypothesis by demonstrating that atomic-fact based retrieval outperforms sub-question based retrieval.
@@
Figure 3: The process of dataset construction and instruction tuning. We first extract the neighbors of the corresponding entity from Wikidata, and then utilize an LLM to convert each triple into a single atomic fact. Subsequently, all facts are transformed into a text comprising multiple molecular clauses by the LLM. The atomic facts are then used to construct each sentence. After two prompts, we obtain the dataset for instruction-tuning with LoRA.
IV Approach
@@
In this section, a novel Atomic fact decomposition-based Retrieval and Editing (ARE) framework is proposed for AQA task. Specifically, our ARE framework contains three stages: 1) Molecular-to-Atomic Fact Decomposition generally decomposes the generated long-form answer into molecular clauses and atomic facts by an instruction-tuned LLM; 2) Atomic Fact based Evidence Retrieval employs a search engine to retrieve evidence and selects the most similar evidence using a reranking model; 3) LLM-based Evidence Verifying and Editing guides LLMs through a verified result to either expand atomic facts for further re-retrieval or edit atomic facts with retrieved evidence.
IV-A Molecular-to-Atomic Fact Decomposition
@@
IV-A1 Long-form answer generation
LLMs are trained on the large-scale corpora which endow them with powerful generative capabilities. Long-form answers are generated by feeding a few-shot prompting template and a specific question to LLMs. Formally, the whole process can be represented as:
@@ IV-A1 Long-form answer generation
IV-A2 Fact decomposition
Fact decomposition aims to decompose long-form answers characterized by complex logical structures into molecular clauses and multiple atomic facts through the instruction-tuning LLM. Specifically, the fact decomposition module consists of two key components: data construction and the instruction-tuning of the Molecule-to-Atomic LLM.
Dataset Construction. To fine-tune the molecule-to-atomic fact decomposition LLM, ARE employs LLMs to construct an instruct-tuning dataset that converts triples into texts containing molecule clauses and atomic facts derived from Knowledge Graphs (KGs) [44, 45]. As depicted in Figure 3, the dataset construction process is divided into two steps: data collection and data processing with LLMs.
Step1: ARE first randomly selects the entities from the Knowledge Graph Question Answering (KGQA) dataset [46] and then extracts the triplets of given entities from Wikidata [^2] https://www.wikidata.org/w/api.php?action=wbsearchentities&search=. As some properties or objects in the triplets may be useless (e.g., JPEG, code, and ISSN), they are thus further filtered by heuristic rules in the extracting process.
Step2: ARE utilizes LLMs to randomly transform several triplets into atomic facts $\{AF_{i1},...,AF_{ik}\}$. The atomic facts are fed into LLMs to generate a response in the Json format: {“Generated content”: $S$, “Molecular clauses”: $MF_{i}$, “Atomic facts”: $AF_{ij}$}, where $S$ is text consisting of $m$ molecular clauses, $MF_{i}$ is $i$-th molecular clause, $i\in\{1,...,m\}$, and $AF_{ij}$ is the $j$-th atomic fact of the $MF_{i}$, $j\in\{1,...,k\}$.To retain the overall coherence and logical flow of the original long form answer, we format the molecular clauses and atomic facts into a Json structure: {“output”:[$MF_{i}$$AF_{ij}$...}…]}, which enhances the relationship among $MF_{i}$, $AF_{ij}$ and $X$.
Through the above process, an instruction-tuning dataset can be obtained, which contains 7138 samples for training and 1000 for evaluation.
Instruction-tuning the Molecule-to-Atomic LLM. For the fine-tuning stage, ARE employs the Low-Rank Adaptation (LoRA) technique [47] for instruction-tuning the Llama3-8B-Instruct LLM to reduce computational complexity and time consumption.
Specifically, LoRA works by freezing parameters $\theta_{0}$ of the pre-trained model, while adding trainable parameters $\Delta\theta_{0}$ that can be expressed as the product of two low-rank matrices:
$\Delta\theta_{0}=\mathbf{BA},$ (6)